# RS-MoE: Mixture of Experts for Remote Sensing Image Captioning and Visual Question Answering

### v1

@@ V-B Extension on the RSVQA Task
To verify the generalization of the proposed model, we further evaluated its performance on the visual question answering (VQA) task for remote sensing images. Specifically, we used the RSEval dataset’s 10 categories of questions, including presence, quantity, color, absolute position, relative position, area comparison, road direction, image, scene, and reasoning. We queried BLIP2, MiniGPT4, InstructBLIP, RSGPT, and our RS-MoE model, all of which were trained on the RSICap dataset, and calculated the average accuracy of their answers for each category. As shown in Table VIII, our model achieved the highest performance in seven out of the ten categories. To visually demonstrate the results on RSVQA, we randomly selected a remote sensing image from the dataset and displayed the answers provided by these models in four aspects: Theme, Quantity, Absolute Position, and Relative Position. It was observed that for simple Theme questions, all models were able to answer accurately. However, as the complexity of the questions increased, especially for those related to spatial relationships, the baseline models’ performance deteriorated, while our model continued to provide accurate answers. This indicates that our model excels in handling complex questions, demonstrating its superior capability in understanding and reasoning about spatial relationships in remote sensing images.
@@ VI Conclusion
This paper introduces RS-MoE, a pioneering Mixture of Experts (MoE)-based Vision-Language Model (VLM) specifically developed for remote sensing image captioning. RS-MoE consists of three core modules: the Image Encoder, VLM Encoder, and MoE Block. At the heart of the framework, the MoE Block leverages an Instruction Router to dynamically generate task-specific prompts, guiding each Large Language Model (LLM) in subsequent layers to focus on distinct aspects of captioning—theme comprehension, object identification, and relationship interpretation. This design allows the model to capture intricate semantic features unique to remote sensing imagery, enabling the generation of detailed, contextually accurate captions. Additionally, we propose a two-stage training strategy to address performance degradation from sparsity issues inherent in MoE-based methods and enhance training efficiency. Our extensive experiments show that, even after fine-tuning on only one dataset, RS-MoE achieves state-of-the-art performance across four RSIC datasets and demonstrates robust generalization capabilities on an RSVQA dataset. Notably, our lightweight RS-MoE-1B variant achieves performance comparable to larger 13B VLMs, underscoring the model’s efficiency in remote sensing applications. Thus, RS-MoE not only excels in performance but also proves practical for real-world applications, particularly in resource-constrained environments where computational efficiency is essential.

### v2

@@ V-B Extension on the RSVQA Task
To verify the generalization of the proposed model, we further evaluated its performance on the visual question answering (VQA) task for remote sensing images. Specifically, we used the RSEval dataset’s 10 categories of questions, including presence, quantity, color, absolute position, relative position, area comparison, road direction, image, scene, and reasoning. We queried BLIP2, MiniGPT4, InstructBLIP, RSGPT, and our RS-MoE model, all of which were trained on the RSICap dataset, and calculated the average accuracy of their answers for each category. As shown in Table X, our model achieved the highest performance in seven out of the ten categories. To visually demonstrate the results on RSVQA, we randomly selected a remote sensing image from the dataset and displayed the answers provided by these models in four aspects: Theme, Quantity, Absolute Position, and Relative Position. It was observed that all models were able to answer simple Theme questions accurately. However, as the complexity of the questions increased, especially for those related to spatial relationships, the baseline models’ performance deteriorated, while our model continued to provide accurate answers. This indicates that our model excels in handling complex questions, demonstrating its superior capability in understanding and reasoning about spatial relationships in remote sensing images.
To evaluate the effectiveness of our model further, we conducted experiments on the RSIVQA dataset, a benchmark for the visual question answering (VQA) task in remote sensing. The RSIVQA dataset includes three question categories: Yes/No, Number, and Others, with the final results summarized in Table XI. Our model is compared against state-of-the-art methods, including MAIN [66], EasyToHard [67], SHRNet [68], and RSAdapter [69]. As shown in the table, our model achieves the best performance across all metrics, with an overall accuracy of 89.12% and an average accuracy of 88.78%. Notably, for the challenging Number category, our model achieves a significant improvement, with an accuracy of 75.46%, outperforming the previous best model, RSAdapter, by 12.82%. Similarly, in the Yes/No and Others categories, our model attains accuracies of 98.01% and 92.88%, respectively, setting new benchmarks in both cases. These results highlight the superior generalization ability of our model in handling diverse and complex question types in the RSIVQA dataset. Its exceptional performance, particularly in the Number and Others categories, underscores its capability to reason about numerical quantities and contextual relationships in remote sensing imagery. This demonstrates the robustness and adaptability of our approach in addressing the unique challenges posed by VQA tasks in remote sensing.
@@ VI Conclusion
This paper introduces RS-MoE, a pioneering Mixture of Experts (MoE)-based Vision-Language Model (VLM) specifically developed for remote sensing image captioning. RS-MoE consists of four core modules: the Image Encoder, VLM Encoder, LLM Block, and MoE Block. At the heart of the framework, the MoE Block leverages an Instruction Router to dynamically generate task-specific prompts, guiding each Large Language Model (LLM) in subsequent layers to focus on distinct aspects of captioning—theme comprehension, object identification, and relationship interpretation. This design allows the model to capture intricate semantic features unique to remote sensing imagery, enabling the generation of detailed, contextually accurate captions. Additionally, we propose a two-stage training strategy to address performance degradation from sparsity issues inherent in MoE-based methods and enhance training efficiency. Our extensive experiments show that, even after fine-tuning on only one dataset, RS-MoE achieves state-of-the-art performance across five RSIC datasets and demonstrates robust generalization capabilities on two RSVQA datasets. Notably, our lightweight RS-MoE-1B variant achieves performance comparable to larger 13B VLMs, underscoring the model’s efficiency in remote sensing applications. Thus, RS-MoE not only excels in performance but also proves practical for real-world applications, particularly in resource-constrained environments where computational efficiency is essential.
While RS-MoE demonstrates significant advantages, there are certain limitations that warrant further exploration. The reliance on predefined sub-tasks for LLM specialization may limit adaptability to highly complex or novel remote sensing scenarios. Additionally, the two-stage training strategy, while effective, introduces additional computational overhead during the pretraining phase. Future work could explore adaptive task routing mechanisms that allow for greater flexibility and generalization to new tasks. Furthermore, integrating multi-modal data sources, such as radar and hyperspectral imagery, could expand the applicability of RS-MoE to a broader range of remote sensing challenges.